# REN: Fast and Efficient Region Encodings from Patch-Based Image Encoders

### v1

@@ Abstract
We introduce the Region Encoder Network (REN), a fast and effective model for generating region-based image representations using point prompts. Recent methods combine class-agnostic segmenters (e.g., SAM) with patch-based image encoders (e.g., DINO) to produce compact and effective region representations, but they suffer from high computational cost due to the segmentation step. REN bypasses this bottleneck using a lightweight module that directly generates region tokens, enabling 60 faster token generation with 35 less memory, while also improving token quality. It uses a few cross-attention blocks that take point prompts as queries and features from a patch-based image encoder as keys and values to produce region tokens that correspond to the prompted objects. We train REN with three popular encoders—DINO, DINOv2, and OpenCLIP—and show that it can be extended to other encoders without dedicated training. We evaluate REN on semantic segmentation and retrieval tasks, where it consistently outperforms the original encoders in both performance and compactness, and matches or exceeds SAM-based region methods while being significantly faster. Notably, REN achieves state-of-the-art results on the challenging Ego4D VQ2D benchmark and outperforms proprietary LMMs on Visual Haystacks’ single-needle challenge. Code and models are available at: https://github.com/savya08/REN.
@@ 1 Introduction
We train REN with multiple image encoders and release the code and trained models to facilitate future research and applications.
@@ 3.1 Training via Self-Supervised Learning and Knowledge Retention
We train REN using images from the Ego4D dataset [13], which offers diverse scenes with objects at varying scales and clutter. For each image encoder, a single REN model is trained and used across all tasks. We preprocess the training dataset to extract segmentation masks using SAM 2 [35]. These masks are used to guide the contrastive and feature similarity losses, as described below.
Contrastive Token Learning. We use the InfoNCE loss [42] to ensure region tokens are consistent across different views of the same object. Specifically, we identify regions in the training images using SAM 2 [35] and assign an ID to each region. Then, we generate two augmented views of each image using random horizontal flipping, rotation, cropping, color jitter, and sharpness adjustments. REN is used to extract region tokens from both views. These "REN tokens" corresponding to prompt points that fall inside the same region (i.e., are fully contained within the same SAM 2-generated mask) are treated as positives, while all others are considered negatives. Specifically, we compute
@@
Feature Similarity. Training only on the contrastive loss could cause the model to drift away from the original encoder’s feature space. This is undesirable, as the original models are well-trained using large and carefully curated datasets, and we would like to retain their generalization capability. To this end, we introduce a feature similarity objective that encourages alignment between region tokens and the original encoder features. Specifically, for each region, we generate a target token by average pooling encoder features within the SAM 2 mask [38], and apply cosine embedding loss against the "aligned REN token" , obtained via a linear projection of the REN token :
@@ 5 Conclusion
Acknowledgements. Work supported in part by ONR N00014-23-1-2383, DARPA HR0011-23-9-0060, NSF grants 2008387, 2045586, 2106825, NIFA award 2020-67021-32799.
@@ Appendix A Implementation Details
Training. Training is performed on images sampled from the Ego4D dataset [13]. We use a batch size of 16 and randomly sample up to 256 point prompts per image to generate region tokens. Segmentation masks are precomputed for all training images using the SAM 2 Automatic Mask Generator [35], with a 3232 point grid and a stability threshold of 0.9. These masks are used to assign region IDs for the contrastive loss and to average-pool encoder features for the feature similarity loss. If multiple masks overlap a given point prompt, we select the mid-sized one.

### v2

@@ Abstract
We introduce the Region Encoder Network (REN), a fast and effective model for generating region-based image representations using point prompts. Recent methods combine class-agnostic segmenters (e.g., SAM) with patch-based image encoders (e.g., DINO) to produce compact and effective region representations, but they suffer from high computational cost due to the segmentation step. REN bypasses this bottleneck using a lightweight module that directly generates region tokens, enabling 60 faster token generation with 35 less memory, while also improving token quality. It uses a few cross-attention blocks that take point prompts as queries and features from a patch-based image encoder as keys and values to produce region tokens that correspond to the prompted objects. We train REN with three popular encoders—DINO, DINOv2, and OpenCLIP—and show that it can be extended to other encoders without dedicated training. We evaluate REN on semantic segmentation and retrieval tasks, where it consistently outperforms the original encoders in both performance and compactness, and matches or exceeds SAM-based region methods while being significantly faster. Notably, REN achieves state-of-the-art results on the challenging Ego4D VQ2D benchmark and outperforms proprietary LMMs on Visual Haystacks’ single-needle challenge. The code and pretrained models are available at https://github.com/savya08/ren.
@@ 1 Introduction
We train REN with multiple image encoders and will release the code and trained models to facilitate future research and applications.
@@ 3.1 Training via Self-Supervised Learning and Knowledge Retention
We train REN using images from the Ego4D dataset [13], which offers diverse scenes with objects at varying scales and clutter. For each image encoder, a single REN model is trained and used across all tasks. We preprocess the training dataset to extract segmentation masks using SAM [22]. These masks are used to guide the contrastive and feature similarity losses, as described below.
Contrastive Token Learning. We use the InfoNCE loss [42] to ensure region tokens are consistent across different views of the same object. Specifically, we identify regions in the training images using SAM [22] and assign an ID to each region. Then, we generate two augmented views of each image using random horizontal flipping, rotation, cropping, color jitter, and sharpness adjustments. REN is used to extract region tokens from both views. These "REN tokens" corresponding to prompt points that fall inside the same region (i.e., are fully contained within the same SAM-generated mask) are treated as positives, while all others are considered negatives. Specifically, we compute
@@
Feature Similarity. Training only on the contrastive loss could cause the model to drift away from the original encoder’s feature space. This is undesirable, as the original models are well-trained using large and carefully curated datasets, and we would like to retain their generalization capability. To this end, we introduce a feature similarity objective that encourages alignment between region tokens and the original encoder features. Specifically, for each region, we generate a target token by average pooling encoder features within the SAM mask [38], and apply cosine embedding loss against the "aligned REN token" , obtained via a linear projection of the REN token :
@@ 4.1 Compute Requirements
| Method | Params (M) | Time (s/img) | |
| SLIC | Grid | ||
| SAM [22] | 641.1 | 2.310 | |
| Batched-SAM [24] | 641.1 | 2.048 | |
| EfficientViT-SAM [51] | 203.3 | 1.790 | |
| SAM 2 [35] | 224.4 | 1.898 | |
[…]
@@ 4.2 Downstream Tasks
| Method | stAP | tAP | Succ. | Rec. |
| SiamRCNN [13] | 0.13 | 0.21 | 41.6 | 34.0 |
| CocoFormer [48] | 0.18 | 0.26 | 48.1 | 43.2 |
| VQLoC [18] | 0.24 | 0.32 | 55.9 | 45.1 |
| HERO-VQL | 0.28 | 0.37 | 60.7 | 45.3 |
| PRVQL [10] | 0.28 | 0.37 | 59.4 | 45.7 |
| RELOCATE [20] | 0.35 | 0.43 | 60.1 | 50.6 |
[…]
@@ 5 Conclusion
Acknowledgements
This work is supported in part by ONR award N00014-23-1-2383, U.S. DARPA ECOLE Program No. #HR00112390060, National Science Foundation grants #2008387, #2045586, #2106825, and NIFA award 2020-67021-32799. We used GPUs at NCSA Delta through allocation CIS240541 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, ONR, or the U.S. Government.
@@ Appendix A Implementation Details
Training. Training is performed on images sampled from the Ego4D dataset [13]. We use a batch size of 16 and randomly sample up to 256 point prompts per image to generate region tokens. Segmentation masks are precomputed for all training images using the SAM Automatic Mask Generator [22], with a 3232 point grid and a stability threshold of 0.9. These masks are used to assign region IDs for the contrastive loss and to average-pool encoder features for the feature similarity loss. If multiple masks overlap a given point prompt, we select the mid-sized one.